OpenStudy (mathhelp346):

(x/2 + 1) - (x/3 - 1) Explain how you solved it step by step please, thank you!!

jhonyy9 (jhonyy9):

x/2 +1 = (x+2)/2 x/3 -1 = (x-3)/3 (x+2)/2 -(x-3)/3 = (3(x-2) -2(x-3))/6 = (3x-6-2x+6)/6 = x/6